Bear population. To estimate the size of the bear population on the Keweenaw Peninsula, conservationists captured, tagged, and released 50 bears. One year later, a random sample of 100 bears included only 2 tagged bears. What is the conservationist’s estimate of the size of the bear population?
Answer by checkley77(12569)

50/x=2/100
2x=50*100
2x=5,000
x=5,000/2
x=2,500 is the estimated number of bears.